# A Study On Mixup-inspired Augmentation Methods For Software Vulnerability Detection

### v1

@@ 4. Results
Finding 1: Stochastic Perturbation performs better than other token-based representation-level methods in both settings. Table 3 and Table 4 show the results of the 5 representation-level augmentation methods when applied blindly (without conditioned augmentation) and conditioned, respectively. We observe that the majority of such augmentation methods lead to performance improvements as they are applied blindly. Binary Interpolation is the only method that degrades the overall F1-score, and Stochastic Perturbation generates the highest gain (9.67%). We observe a similar trend in conditioned augmentation variants, where all methods achieved a performance gain compared with no augmentation, and Stochastic Perturbation performs the best among the conditioned representation-level augmentation approaches as well and achieves an improvement (8.92%) over No Augmentation in terms of F1-score.
Finding 3: All representation-level augmentation methods beat VGX, a SOTA vulnerability generation method. Intriguingly, VGX (A SOTA vulnerability generation method that outperforms Vulgen) fails to increase performance when used for balancing a real-world dataset like BigVul, which is more imbalanced and bigger than other datasets, such as Devign (Zhou et al., 2019) and Reveal (Chakraborty et al., 2022).
Table 3. Blind Augmentation Results. The ones beating No Augmentation are marked in blue, otherwise, marked in red.
@@
Table 4. Conditioned Augmentation Results. The ones beating No Augmentation are marked in blue, otherwise, marked in red.
@@ 5.1. Why does Stochastic Perturbation’s performance degrade when conditioned?
As we delve into the mechanism of Stochastic Perturbation, we realize that using this method to augment data is equivalent to using a DropOut layer in the network, which gets omitted one out of 23 times (the original data) it sees a vulnerable sample. However, in the conditioned setting, since the vulnerable tokens will be reinstated after the DropOut, they will behave differently from the original DropOut’s regularization effect and the expected value of the output will no longer be equal to that of the input data. This is while other methods do not have such an effect in the naive setting, and keeping the vulnerable section proves to be useful for them.
@@ 5.4. Implications of our findings
We recommend that future researchers consider using Random Oversampling instead of token-bsaed representation-level vulnerability augmentation when aiming for a balanced dataset. However, one might prefer to keep the original ratio and augment the minority class while adding new items of the majority class that are widely available. In such a case, Stochastic Perturbation seems the best augmenting strategy to create new items. This is because adding new clean items while repeating the same vulnerable samples will only add the unwanted noise of the new clean items, without adding any extra useful information about the vulnerabilities. Also, based on the results of VulScriber (Daneshvar et al., 2024) and ours, we can conclude that VGX (Nong et al., 2024) is not a practical method for dealing with data shortage when we aim for either balancing a dataset or keeping the original ratio.
Lastly, our examination of conditioned methods reveals that while they offer potential, they still introduce a notable amount of noise. Given these observations, we advocate for future investigations to delve deeper into more intrusive augmentation techniques for DLVD. This recommendation contrasts with prior studies that found success with similar methodologies for different SE tasks.
@@ 5.5. Threats to Validity And Limitations
Additionally, a threat concerns the generalization capacity of our method, influenced by the choice of the model. Since our study focuses on augmentation, and many algorithms are tested concurrently, our analysis does not cover a wide range of Vulnerability Detection models. We exclusively employed LineVul, a recent DLVD model, due to its high performance, partly influenced by time constraints and resource availability. In the future, we plan to include other models to validate the effect of our augmentation methods.
6. Conclusion
We conducted the first comprehensive study on the impact of representation-level augmentation in the context of Deep Learning-based Vulnerability Detection. Our study introduced a vulnerability detection-specific approach to augmenting data, conditioning the augmentation based on the vulnerable segment in the embedding. Our findings offer valuable insights and recommendations for practitioners in this field. Notably, Stochastic Perturbation emerged as the top-performing method. We also showed that conditioning the augmentation is beneficial for most representation-level augmentation methods.
For Future work, we plan to investigate the effect of graph-based augmentation methods (e.g. Manifold Mixup) that aside from the text representation, also take the graph representation of the code (e.g. AST or CFG) into account.
7. Data Availability
Our code is available here for replication and further improvements.

### v2

@@ 4. Results
Table 3. Blind augmentation results. The ones beating No augmentation are marked in blue; otherwise, marked in red.
@@
Table 4. Conditioned augmentation results. The ones beating No augmentation are marked in blue; otherwise, marked in red.
@@
Finding 1: Stochastic Perturbation performs better than other token-based representation-level methods in both settings. Tables 3 and 4 show the results of the five representation-level augmentation methods when applied blindly (without conditioned augmentation) and conditioned, respectively. We observe that the majority of such augmentation methods lead to performance improvements as they are applied blindly. Binary Interpolation is the only method that degrades the overall F1-score, and Stochastic Perturbation generates the highest gain (9.67%). We observe a similar trend in conditioned augmentation variants, where all methods achieved a performance gain compared with no augmentation, and Stochastic Perturbation performs the best among the conditioned representation-level augmentation approaches as well and achieves an improvement (8.92%) over No Augmentation in terms of F1-score.
Finding 3: All representation-level augmentation methods beat VGX, a SOTA vulnerability generation method. Intriguingly, VGX (a SOTA vulnerability generation method that outperforms Vulgen) fails to increase performance when used for balancing a real-world dataset like BigVul, which is more imbalanced and bigger than other datasets such as Devign (Zhou et al., 2019) and Reveal (Chakraborty et al., 2022).
@@ 5.1. Why Does Stochastic Perturbation’s Performance Degrade When Conditioned?
As we delve into the mechanism of Stochastic Perturbation, we realize that using this method to augment data is equivalent to using a DropOut layer in the network, which gets omitted one out of 23 times (the original data) it sees a vulnerable sample. However, in the conditioned setting, since the vulnerable tokens will be reinstated after the DropOut, they will behave differently from the original DropOut’s regularization effect, and the expected value of the output will no longer be equal to that of the input data. This is while other methods do not have such an effect in the naive setting, and keeping the vulnerable section proves to be useful for them.
@@ 5.4. Implications of Our Findings
We recommend that future researchers consider using Random Oversampling instead of token-based representation-level vulnerability augmentation when aiming for a balanced dataset. However, one might prefer to keep the original ratio and augment the minority class while adding new items of the majority class that are widely available. In such a case, Stochastic Perturbation seems the best augmenting strategy to create new items. This is because adding new clean items while repeating the same vulnerable samples will only add the unwanted noise of the new clean items, without adding any extra useful information about the vulnerabilities. Also, based on the results of VulScriber (Daneshvar et al., 2024) and ours, we conclude that VGX (Nong et al., 2024) is not a practical method for dealing with data shortage when we aim for either balancing a dataset or keeping the original ratio.
Lastly, our examination of conditioned methods reveals that while they offer potential, they still introduce a notable amount of noise. Given these observations, we advocate for future investigations to delve deeper into more intrusive augmentation techniques for DLVD. This recommendation contrasts with prior studies that found success with similar methodologies for different software engineering tasks.
@@ 5.5. Threats to Validity and Limitations
Additionally, a threat concerns the generalization capacity of our method, which is influenced by the choice of the model. Since our study focuses on augmentation, and many algorithms are tested concurrently, our analysis does not cover a wide range of Vulnerability Detection models. We exclusively employed LineVul (a recent DLVD model) due to its high performance, partly influenced by time constraints and resource availability. In the future, we plan to include other models to validate the effect of our augmentation methods.
6. Conclusions
In this paper, we conducted the first comprehensive study on the impact of representation-level augmentation in the context of Deep Learning-based Vulnerability Detection. Our study introduced a vulnerability detection-specific approach to augmenting data, conditioning the augmentation based on the vulnerable segment in the embedding. Our findings offer valuable insights and recommendations for practitioners in this field. Notably, Stochastic Perturbation emerged as the top-performing method. We also showed that conditioning the augmentation is beneficial for most representation-level augmentation methods.
As for future work, we plan to investigate the effect of graph-based augmentation methods (e.g., Manifold Mixup) that, aside from the text representation, also take the graph representation of the code (e.g., AST or CFG) into account.
7. Code Availability
Our code is available at https://doi.org/10.5281/zenodo.13916933 for replication and further improvements.
Acknowledgments
This work is partially supported by NSERC (Canada) and the University of Manitoba.